v6mpiresk8ter:

You are a volunteer for the school store. One of the most popular items is strawberry-banana-orange juice. There are two local vendors that will deliver the juice to the school store at the beginning of every month. Healthy Drinks, Inc. charges $0.39 per bottle with a delivery fee of $25. The Squeeze charges $0.18 per bottle with a delivery fee of $65. a. You want to stock the store with juice at the beginning of the school year. How much will it cost to purchase 300 bottles of strawberry-banana-orange juice from Healthy Drinks, Inc.?

v6mpiresk8ter:

$142 $92 $325.39 $117 (question choices)

surjithayer:

any idea?

IhelpUuHELPme:

o-o Well as you can see- Healthy Drinks Inc. charges 0.39 PER bottle with a delivery fee of 25 Your equation would be: \[0.39b + 25 = total\] b represents bottles (: As you see, the amount of bottles is already given to us, which is 300, so you replace b with 300! \[0.39(300) +25 = total\] When there's no fraction and there's just a number in brackets next a another number, you multiply (: so \[0.39 \times 300\] which is 117 The delivery fee is $25 so you add 25 to 117 which is 142 (: So it would cost $142 to buy 300 bottles of strawberry-banana-orange juice from Healthy Drinks Inc. <3 -We don't need to worry bout the other vender because the question didn't mention it

Zari:

add all the numbers up and see what it come up too